If the one-year spot rate on a security is 8 percent (YTM) and the one-year forward rate is 12.04%, what should the spot rate of the corresponding security with a two year maturity?

Group of answer choices 4.04% 0.40% 16.04% 8.02% 10.00%
Answer #1

Using Pure Expectation Theory,

(1 + r1,2)2 = (1 + r1,1)(1 + r2,1)

(1 + r1,2)2 = (1.08)(1.1204)

r1,2 = 10.00%